# Energy Consumption of TLS, Searchable Encryption and Fully Homomorphic Encryption

### v2

@@ Abstract
This work investigates this question by measuring the energy consumption increase induced by three PETs compared to their non-private counterparts: TLS, Searchable Encryption, and Fully Homomorphic Encryption (FHE). These technologies were chosen for two reasons. First, they cover different maturity levels –from the widely deployed TLS protocol to the emerging FHE schemes– allowing us to examine the influence of maturity on energy consumption. Second, they each have well-established applications in industry: web browsing, encrypted databases, and privacy-preserving machine learning.
Our results reveal highly variable energy consumption increases, ranging from $2\times$ for TLS to $10\times$ for Searchable Encryption and $100{,}000\times$ for FHE. Our experiments demonstrate a simple and reproducible methodology, based on existing open-source software, to quantify the energy costs of PETs. They also highlight the wide spectrum of energy demands across technologies, underscoring the importance of further research on sustainable PET design. Finally, we discuss orthogonal research directions, such as hardware acceleration, to outline promising directions toward sustainable PETs.
@@ 2.1 Life-Cycle Assessment
One may wonder why the hardware is not always included in the setup. Even though hardware is a significant part of the ICT footprint [17], the hardware-related costs are irrelevant in the scope of these works. These works [19, 35] (like ours) want to identify the most energy-efficient services and execute them on a standard hardware. In other words, the hardware-related footprint is excluded because they want to minimize the footprint of the service usage.
@@ 3.1 Experimental setup
As the hardware can slightly change the energy consumption of an algorithm, one may wonder whether we should execute our experiment on different devices. However, remember that we mainly want to measure the relative footprint difference between private and non-private deployments. A standard hardware change would influence the absolute values, but the relative difference should not be significantly changed. This measurement on a single hardware platform aligns with common practices in similar studies on ML algorithms [32, 15].
@@ 3.2 Fully Homomorphic Encryption (FHE)
FHE is a type of cryptography scheme [7, 8] allowing to perform arithmetic operations on encrypted data. With the increasing use of ML, there is a growing concern regarding the use of personal data in these systems. To address this problem, FHE has appeared as a nice solution to execute traditional ML pipelines under encryption.
@@ 3.2.1 Encrypted Machine Learning inference
To better understand this overhead, Figure 2(a) studies the influence of the number of features on the footprint of three popular classification models: logistic regression, random forest, and neural networks. We observe that the scaling of the encrypted inferences does not match perfectly the scaling of the plaintext inferences. First, the encrypted logistic regression footprint increases relatively faster than the cost of the plaintext inference. Second, the encrypted neural network seems to have the same scaling pattern as the encrypted logistic regression. The encrypted neural network does not have results for more than 200 features because we stopped inferences taking more than several hours (to limit the energy consumption of our experiments). Third, the encrypted random forest footprint is approximately constant. This behavior seems to be an artifact coming from the fact that Concrete ML transforms decision trees into matrix multiplications (inducing significant constant costs).
Figure 2(b) illustrates the footprint in function of the number of samples. On the one hand, we see that plaintext algorithms are optimized to process batches of data points: the average cost is amortized when the number of samples grows. On the other hand, we observe on the logistic regression that the encrypted inference does not provide a similar amortization. Like on Figure 2(a), we stopped experiments requiring several hours (which explains the smaller number of results for the random forest and neural network).
@@ 3.3 Searchable Encryption
Results. Figure 4 shows the results of this benchmark for varying database sizes. We observe that the footprint of the encrypted database is nearly $10\times$ higher than the footprint of the plaintext database. Moreover, the footprint of the encrypted database increases slightly faster in function of the database size.
@@ 3.4 TLS
Unfortunately, we cannot compare our results to existing works on the energy consumption of TLS [29, 31]. While we analyze the energy consumption to estimate the environmental footprint of HTTPS, these works had a completely different goal: estimating the impact of HTTPS on a smartphone battery life. Thus, these works have only measured client-side energy consumption.
@@
Discussions. Our experiments only simulate static websites, and exclude any possible back-end operations (e.g., database interactions), even though such operations are common in Web applications. To assess the footprint of a Web application, an experiment would include both the cost related to HTTP(S) and the back-end costs. However, this is not our goal: we only want to compare HTTP to HTTPS. Hence, we exclude any backend work to focus solely on the footprint of HTTP data transfer.
@@ 5 Conclusion
However, this correlation should not hide the fact that FHE remains several orders of magnitude more expensive than its plaintext equivalent. While performance optimization is a central goal in cryptography research, this raises a key question: what is the optimal performance? Modern FHE cryptosystems rely on inherently expensive primitives, unlike TLS, which makes reaching comparable efficiency difficult, if not impossible. This observation should not discourage research on FHE, but it illustrates a potential broader debate for the cryptography community about the trade-off between privacy and environmental footprint.
@@ 5.0.1 Acknowledgements
This work was supported by the Netherlands Organization for Scientific Research (De Nederlandse Organisatie voor Wetenschappelijk Onderzoek) under NWO:SHARE project [CS.011].
A Large Language Model has been carefully used to reformulate some sentences and fix typos.

### v3

@@ Abstract
This work investigates this question by measuring the energy consumption increase induced by three PETs compared to their non-private equivalents: TLS, Searchable Encryption, and Fully Homomorphic Encryption (FHE). These technologies were chosen for two reasons. First, they cover different maturity levels –from the widely deployed TLS protocol to the emerging FHE schemes– allowing us to examine the influence of maturity on energy consumption. Second, they each have well-established applications in industry: web browsing, encrypted databases, and privacy-preserving machine learning.
Our results reveal highly variable energy consumption increases, ranging from $2\times$ for TLS to $10\times$ for Searchable Encryption and $100{,}000\times$ for FHE. Our experiments demonstrate a simple and reproducible methodology, based on existing open-source software, to quantify the energy consumption of PETs. They also highlight the wide spectrum of energy demands across technologies, underscoring the importance of further research on sustainable PET design. Finally, we discuss orthogonal research directions, such as hardware acceleration, to outline promising directions toward sustainable PETs.
@@ 2.1 Life-Cycle Assessment
One may wonder why the hardware is not always included in the setup. Even though hardware is a significant part of the ICT footprint [17], the hardware-related costs are irrelevant in the scope of these works. These works [19, 35] (like ours) want to identify the most energy-efficient services and execute them on a standard hardware. In other words, the hardware-related footprint is excluded because they want to minimize the consumption of the service usage.
@@ 3.1 Experimental setup
As the hardware can slightly change the energy consumption of an algorithm, one may wonder whether we should execute our experiment on different devices. However, remember that we mainly want to measure the relative consumption difference between private and non-private deployments. A standard hardware change would influence the absolute values, but the relative difference should not be significantly changed. This measurement on a single hardware platform aligns with common practices in similar studies on ML algorithms [32, 15].
@@ 3.2 Fully Homomorphic Encryption (FHE)
FHE is a type of cryptographic scheme [7, 8] allowing to perform arithmetic operations on encrypted data. With the increasing use of ML, there is a growing concern regarding the use of personal data. To address this problem, FHE has appeared as a nice solution to execute traditional ML pipelines under encryption.
@@ 3.2.1 Encrypted Machine Learning inference
To better understand this overhead, Figure 2(a) studies the influence of the number of features on the consumption of three popular classification models: logistic regression, random forest, and neural networks. We observe that the scaling of the encrypted inferences does not match perfectly the scaling of the plaintext inferences. First, the encrypted logistic regression consumption increases relatively faster than the cost of the plaintext inference. Second, the encrypted neural network seems to have the same scaling pattern as the encrypted logistic regression. The encrypted neural network does not have results for more than 200 features because we stopped inferences taking more than several hours (to limit the energy consumption of our experiments). Third, the encrypted random forest consumption is approximately constant. This behavior seems to be an artifact coming from the fact that Concrete ML transforms decision trees into matrix multiplications (inducing significant constant costs).
Figure 2(b) illustrates the consumption in function of the number of samples. On the one hand, we see that plaintext algorithms are optimized to process batches of data points: the average cost is amortized when the number of samples grows. On the other hand, we observe on the logistic regression that the encrypted inference does not provide a similar amortization. Like on Figure 2(a), we stopped experiments requiring several hours (which explains the smaller number of results for the random forest and neural network).
@@ 3.3 Searchable Encryption
Results. Figure 4 shows the results of this benchmark for varying database sizes. We observe that the consumption of the encrypted database is nearly $10\times$ higher than the consumption of the plaintext database. Moreover, the consumption of the encrypted database increases slightly faster in function of the database size.
@@ 3.4 TLS
Unfortunately, we cannot compare our results to existing works on the energy consumption of TLS [29, 31]. While we analyze the energy consumption of the whole HTTPS protocol, these works had a completely different goal: estimating the impact of HTTPS on a smartphone battery life. Thus, these works have only measured client-side energy consumption.
@@
Discussions. Our experiments only simulate static websites, and exclude any possible back-end operations (e.g., database interactions), even though such operations are common in Web applications. To assess the footprint of a Web application, an experiment would include both the cost related to HTTP(S) and the back-end costs. However, this is not our goal: we only want to compare HTTP to HTTPS. Hence, we exclude any backend work to focus solely on the consumption of HTTP data transfer.
@@ 5 Conclusion
However, this correlation should not hide the fact that FHE remains several orders of magnitude more expensive than its plaintext equivalent. While performance optimization is a central goal in cryptography research, this raises a key question: what is the optimal performance? Modern FHE schemes rely on inherently expensive primitives, unlike TLS, which makes reaching comparable efficiency difficult, if not impossible. This observation should not discourage research on FHE, but it illustrates a potential broader debate for the cryptography community about the trade-off between privacy and environmental footprint.
@@ 5.0.1 Acknowledgements
This work was supported by the Netherlands Organization for Scientific Research (De Nederlandse Organisatie voor Wetenschappelijk Onderzoek) under NWO:SHARE project [CS.011]. We thank Ségolène Blanchard for proofreading. A Large Language Model was carefully used to polish some paragraphs.